OpenStudy (anonymous):

Annuity question. Steve deposits $50.00 each week into a bank account that earns 5.75% interest, compounded weekly. How much money will he have saved after 90 weeks? @AravindG

OpenStudy (campbell_st):

well the future value of an annuity is \[A = \frac{M[(1 + r)^n -1]}{r}\] M is the contribution paid at the end of the period. r is the interest rate as a decimal n = number of time periods.

OpenStudy (campbell_st):

so the interest is per annum... you need to convert it to weekly then change it to a decimal. The formula above is for a annuity paid at the end of the period.

OpenStudy (anonymous):

convert the equation ?

OpenStudy (campbell_st):

well you need to find the interest and find out when the regular payment is made... start or end of the week as it affects the future value.

OpenStudy (anonymous):

50[(1+0.0575)90 -1] / r 50[(1.0575)^90-1]/r ? ... can't you tell I'm awful with math

OpenStudy (campbell_st):

you forgot 0.575 in the denominator... and you'll get an answer... but the question is extremely poor with information.

OpenStudy (anonymous):

is it 0.575 or 0.0575

OpenStudy (campbell_st):

5.75% = 0.0575

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

N = total number of payments I = annual interest rate (as percent, NOT decimal) PV = present value PMT = regular payment amount FV = future value P/Y = payment periods per year (annual =1, monthly = 12, quarterly = 4, etc.) C/Y = compounding periods per year PMT: this is always END for ordinary annuities A=P[(1+r)n−1]/r A= 50[(1.0575)^90-1]/0.0575 A=50(152.19)/0.0575 A=7609.56/0.0575 ????? is this right so far or am I completely wrong

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

someone help... please

OpenStudy (amistre64):

how much does he have in his account at the end of the first week if he deposits $50 into his account at the end of the week??

OpenStudy (amistre64):

.... this is a method to determine a recurrsion that develops the formula.

OpenStudy (amistre64):

0 week1 0k 50 |---------------|----| = 50 50 week2 50k 50 |---------------|----| = 50k + 50 50k + 50 week3 (50k+50)k 50 |--------------------|--------| = 50k^2 + 50k + 50 i see a pattern forming such that: for the nth week, he will have: 50k^(n-1) + 50k^(n-2) + ... + 50 in the account we can factor out the 50 and the rest becomes a geometric sum 50(1 + k^2 + k^3 + ... + k^(n-1))

OpenStudy (amistre64):

a geometric sum can be proofed out into the formula: (1-k^n)/(1-k)

OpenStudy (amistre64):

the value for k is the compounding factor (1+r/52) in this case